OpenStudy (anonymous):

Let f(X) be the square of the distance from the oint (2,1) to a point (x, 3x+2) on the line y=3x+2. Show that f(x) is a quadratic function, and find its minimum value by completing the square.

ganeshie8 (ganeshie8):

f(x) = square of distance between points (2, 1) and (x, 3x+2)

ganeshie8 (ganeshie8):

use the distance formula and write down f(x)

OpenStudy (anonymous):

is that it?

ganeshie8 (ganeshie8):

wat did u get after applying distance formula ?

OpenStudy (anonymous):

I got root (-10x^2 +13)

ganeshie8 (ganeshie8):

ok, lets see if its correct :)

ganeshie8 (ganeshie8):

we have the distance formula , distance between two points = \(\large \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\)

ganeshie8 (ganeshie8):

distance between points (2, 1) and (x, 3x+2) = \(\sqrt{(x-2)^2+(3x+2-1)^2}\) \(\sqrt{(x-2)^2+(3x+1)^2}\) \(\sqrt{x^2+4-4x+ 9x^2+1+6x}\) \(\sqrt{10x^2+2x+5}\)

ganeshie8 (ganeshie8):

f(x) = square of distance between points (2, 1) and (x, 3x+2)

ganeshie8 (ganeshie8):

so, f(x) = \((\sqrt{10x^2+2x+5})^2\) = \(10x^2+2x+5\)

ganeshie8 (ganeshie8):

thats our required function f(x), clearly, its a quadratic (why ?)

OpenStudy (anonymous):

Oh I see, I foiled it wrong thank you so much!

ganeshie8 (ganeshie8):

np :) you still need to find its minimum value...

OpenStudy (anonymous):

Its minimum value would be the lowest point on the graph c:

ganeshie8 (ganeshie8):

Yes, u knw how to find it ha ? :)

OpenStudy (anonymous):

Can you review over how to for me?

ganeshie8 (ganeshie8):

ok sure :) we simply need to change the quadratic to form : \(\large a(x-h)^2+\color{red}{k}\) \(\color{red}{k}\) is our required minimum value

ganeshie8 (ganeshie8):

\(10x^2+2x+5\) factor out 10 \(10(x^2+2/10x+5/10)\)

OpenStudy (anonymous):

oh okay thank you! I know now, 49/10

ganeshie8 (ganeshie8):

\(10x^2+2x+5\) factor out 10 \(10(x^2+2/10x+5/10)\) \(10(x^2+2(1/10)x+ 1/100 - 1/100 + 5/10)\) \(10((x+1/10)^2 - 1/100 + 5/10)\) \(10((x+1/10)^2 - 1/100 + 50/100)\) \(10((x+1/10)^2 + 49/100)\) \(10(x+1/10)^2 + 10 \times 49/100\) \(10(x+1/10)^2 + 49/10\)

ganeshie8 (ganeshie8):

Yes, \(\color{red}{k}\) = 49/10, so thats our min value !